OpenStudy (anonymous):

What is the x-intercept of the graph y = -4x + 8? -4 0 2 8 I've done my work and came out with 2, but I'm still a bite unsure.

jimthompson5910 (jim_thompson5910):

2 is correct

jimthompson5910 (jim_thompson5910):

check: y = -4x + 8 y = -4(2) + 8 y = -8 + 8 y = 0 so that confirms the answer

OpenStudy (anonymous):

Thanks!

jimthompson5910 (jim_thompson5910):

yw